OpenStudy (anonymous):

Solve for x. |2x + 2| = 14 Solve for x. |2x + 6| - 4 = 20

OpenStudy (anonymous):

Just do it the same way i just told you. When removing the absoulte value you get two cases. 2x+2=14 and 2x+2=-14

OpenStudy (anonymous):

What you want to do with the second one is to isolate the absolute value. So just add 4 to each side and solve ( |2x+6| = 24 ) it like the first one.

OpenStudy (amorfide):

|2x + 2| = 14 |x| means that it is always going to be a positive outcome next look at 14 you know that |14| = 14 but you also know that |-14| = 14 so you know that inside the || you have |2x+2| you know that you must get 14 or -14 inside the || once you get 14 or -14 you know that both answers will turn out to be 14 so this means you will get 2 values of x so since 2x+2 can be 14 or -14 you can write two equations 2x+2=14 2x+2=-14 now solve for x

OpenStudy (anonymous):

how do i do that?

OpenStudy (amorfide):

isolate x you have 2x+2=14 and 2x+2=-14 lets look at 2x+2=14 you want all x terms on one side and all constants on another side so you want to get rid of the +2 on the left side how would you do this?

OpenStudy (anonymous):

minus 2 on both sides

OpenStudy (amorfide):

yes this gives 2x=12 now you want x on it's own you have 2x so what do you do to get x

OpenStudy (anonymous):

divide by 2

OpenStudy (amorfide):

and what do you get?

OpenStudy (anonymous):

6

OpenStudy (amorfide):

yes so your first answer is x=6 now look at 2x+2=-14 what must you do to get rid of the +2 then what must you do to get x what do you get?

OpenStudy (anonymous):

i got -8

OpenStudy (amorfide):

well done so your answer is x=6 or x=-8

OpenStudy (anonymous):

thanks